OpenStudy (anonymous):

Solve -11x = 22 for x. Solve 3x - 4 = 11 for x.

OpenStudy (blank ):

-11x=22 Divide each term in the equation by -11. -(11x)/(-11)=(22)/(-11) Simplify the left-hand side of the equation by canceling the common terms. x=(22)/(-11) Simplify the right-hand side of the equation by simplifying each term. x=-2 Hope this helps. :)

OpenStudy (anonymous):

-2